OpenStudy (anonymous):

Good gracious this is irritating...Divide and simplify to form a+bi; 6-5i over 7+4i???:(

OpenStudy (anonymous):

\[\frac{22}{65}-\frac{59}{65}i\] you want the method?

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

then cheat i will show you how

OpenStudy (anonymous):

\[\frac{6-5i}{7+4i}\] Multiply the top and the bottom by the conjugate of the denominator.\[\frac{6-5i}{7+4i} * \frac{7-4i}{7-4i}\]\[\frac{(6-5i)(7-4i)}{49 - 16i^2}\] \[\frac{(6-5i)(7-4i)}{65}\]\[\frac{42 - 24i - 35i + 20i^2}{65}\] \[\frac{22 - 59i}{65}\]\[\frac{22}{65} - \frac{59}{65}i\]

OpenStudy (anonymous):

thanks tyler!

OpenStudy (anonymous):

that is the method all right. since this is on line work, you can also use a calculator.